# Structure-Fair Quantum Circuit Complexity: An Auditable Information-Theoretic Lower Bound

### v2

@@ Abstract
Keywords Quantum circuit complexity , Structural fairness , Quantum benchmarking , Single-shot quantum information theory , Relative entropy , Kolmogorov complexity
1 Introduction
@@
What, precisely, might constitute the intrinsic complexity of a many-body quantum state? Driven by quantum information science, this fundamental question is increasingly becoming a bridge connecting computational theory and fundamental physics, profoundly influencing our understanding of the potential of quantum computation, the black hole information paradox [1, 2] and its later developments [3], the emergence of spacetime geometry [4, 5], and even phases of matter. At the core of answering this question lies the quantification of the minimum computational resources required to prepare a specific quantum state. This perspective is not only key to assessing the potential of quantum algorithms but also provides a fundamental benchmark for measuring the true performance of quantum hardware, with concrete diagnostics developed in recent works [6, 7]. Although various profound paradigms for complexity measures have been developed in the field—including Nielsen’s complexity based on geometric shortest paths[8], Krylov complexity based on operator algebra growth [9], and resource theories closely related to quantum computational advantage (e.g., magic-state and related approaches [10, 11], recent links to operator growth [12], and general resource-theoretic formalisms [13])—a long-standing challenge remains to be systematically addressed: how to conduct a fair complexity accounting for a “structured” quantum system. The so-called “structural fairness problem” originates from the ubiquitous presence of encodings and symmetries in quantum information processing, which can dramatically reshape what should count as a genuine computational cost [8]. Taking a fault-tolerant quantum computer as an example, its information is encoded within a certain “code subspace” of a vast physical Hilbert space; for a basic operation executed on a logical qubit, its true computational cost should depend on the change in the logical information, rather than the apparent consumption of mobilizing many physical qubits to maintain the error-correcting code structure [14, 15]. In practice, logical operations may be realized via nontrivial transformations supported by the code architecture itself [16, 17]. If a complexity measure cannot distinguish between these two, it could lead to the misleading conclusion that the larger the physical scale of the system and the more complex the encoding structure, the higher its complexity is assessed to be [13]. Similarly, in symmetry-protected topological (SPT) phases or strongly correlated systems, the low-energy physics is confined to specific symmetry sectors; a fair measure ought to treat these structures—endowed “for free” by the Hamiltonian’s symmetries—as an a priori background, and only measure the cost of generating non-trivial quantum correlations on top of this background [18, 19]. This viewpoint is consistent with broader perspectives on emergent phases and constrained dynamics [20]. Existing frameworks for measures often find it difficult to systematically isolate these “free structures,” leading to a lack of a unified and fair benchmark when making cross-comparisons of complexity across different physical systems or encoding schemes [10, 11]. To directly address this challenge, this paper establishes a theory for quantum complexity lower bounds that is structurally fair, theoretically rigorous, and operationally accessible. Its core is the introduction and systematic construction of the theoretical framework of “Reference-Contingent Complexity” (RCC), and its core contributions can be summarized into the following three points:
@@
The structure of this paper is organized as follows. We first systematically establish the theoretical framework for Reference-Contingent Complexity (RCC) in Sec. 2, and then proceed in Sec. 3 to propose and prove the central theorem of its role as a strict lower bound for circuit complexity. To connect this theory with physical reality, Sec. 4 delves into the operational principles of RCC. Building on this foundation, Sec. 5 presents the axiomatic properties of RCC and its three major application interfaces, while Sec. 6 discusses it within a broader theoretical context. Sec. 7 provides a summary of the entire paper. To ensure the self-consistency of the manuscript and the completeness of the arguments, all technical derivations are systematically organized in the appendices. Among them, the core theoretical cornerstones—the rigorous proof of the main theorem, the robustness of the framework, and the operational protocols—are elaborated in detail in Appendices A, B, C, and D, respectively. Building upon this, we further extend the boundaries of the theory: Appendices E and F explore the connections of this framework with other theories and its generalization to broader settings, while Appendix G provides constructive evidence for the central conjecture in the main text. Finally, to support the key application interfaces in the main text, Appendix I provides a complete technical supplement on its connection to thermodynamics. A quick reference list of core symbols and terminology can be found in Appendix H.
2 Theoretical Framework: Reference-Contingent Complexity (RCC)
@@
2.1 System Setting and Notational Conventions
We consider a finite-dimensional quantum system defined on a Hilbert space $H$ of total dimension $d$. The dynamical evolution of the system is driven by a universal finite gate set $G\subseteq\mathrm{SU}(d)$ with cardinality $g$. Our fundamental objective is to quantify the minimum resource cost required to prepare a target state $\rho$ from a simple initial state (such as a product state) using a quantum circuit composed of this gate set. This minimum resource cost, the optimal circuit complexity $C^{*}_{\mathrm{opt}}(\epsilon)(\rho)$, is defined as the length of the shortest sequence of atomic instructions capable of preparing $\rho$ within a precision $\epsilon$ (as measured by the trace distance). To ensure that our complexity benchmark can be compared with concrete physical implementations, we must objectively and unambiguously calibrate the fundamental unit of computational operations. This leads to the core postulate of our framework:
@@ 2.1 System Setting and Notational Conventions
The value of $\Gamma_{R}$ is based on a rigorously formalized computational model that combines the gate set $g$ with the hardware’s addressability $|S_{R}|$. A complete exposition of this model, including the precise definition of $\Gamma_{R}=g\cdot|S_{R}|$ and the handling of parallel circuits and continuous-parameter gates, is axiomatically defined in detail in Appendix A. This postulate directly links abstract computational steps with the physical capabilities of the hardware (gate set size and addressing range), enabling our theory to capture the essential differences in computational power among different physical platforms. To ensure the self-consistency of the theory, we list here the key information-theoretic quantities that will be used in the subsequent derivations. The von Neumann entropy of a quantum state $\rho$ is defined as $S(\rho)=-\mathrm{Tr}(\rho\cdot\log(\rho))$[21]. The quantum relative entropy between two quantum states $\rho$ and $\sigma$ is $D(\rho||\sigma)=\mathrm{Tr}[\rho\cdot(\log(\rho)-\log(\sigma))]$;[22] on the other hand, throughout this paper, to maintain consistency with information-theoretic conventions, all logarithms are assumed to be base-2 (units in bits) by default. In Sec. 5.4 and Appendix I, where we interface with the framework of thermodynamics, we will primarily use the natural logarithm (units in nats) for convenience of calculation, and will make explicit conversions when necessary.
@@
2.2 Reference Set and Structured Priors
@@
2.3 Definition of Reference-Contingent Complexity
@@
$C_{R}(\rho)=\frac{1}{\log\Gamma_{R}}D(\rho||\sigma_{R})$ (2.2)
@@
Eq. (2.3) reveals the physical meaning of RCC. Its numerator term, $(\log d_{R}-S(\rho))$, represents the “information deficit” or “degree of order” of the quantum state $\rho$ within its respective structured subspace $H_{R}$. $\log d_{R}$ is the maximum Shannon information capacity of this subspace (i.e., $S(\sigma_{R})$), while $S(\rho)$ is the von Neumann entropy actually contained in the state $\rho$. The difference between the two is precisely the degree to which the state deviates from “complete disorder,” and it is also precisely the amount of non-trivial information that must be “injected” by any quantum circuit that prepares the state. At the same time, the denominator term $\log\Gamma_{R}$ provides a precise calibration of the fundamental unit of computational resources. Therefore, the physical picture of RCC is: the structurally fair complexity of a quantum state is equal to the minimum amount of information required to be injected for its preparation, divided by the maximum information bandwidth that a single complete computational operation can provide. To specifically illustrate how RCC achieves structural fairness, we once again consider a stabilizer code that encodes $k$ logical qubits into $N$ physical qubits. The dimension of its code subspace $H_{R}$ is $d_{R}=2^{k}$. For a logical pure state $\rho_{L}$ within the code space, its von Neumann entropy is $S(\rho_{L})=0$. Assuming logical gates can act arbitrarily on the $k$ logical qubits with a universal gate set of cardinality $g_{L}$, and that a single operation has an addressable range covering all $k$ logical qubits, i.e., $|S_{R}|=k$. According to Eq. (2.3) and the definition $\Gamma_{R}=g_{L}\cdot|S_{R}|$, its RCC is $C_{R}(\rho_{L})=k/\log(g_{L}\cdot k)$. This result aligns with physical intuition: the complexity of this logical pure state is proportional to the number of logical information bits it carries, $k$, and is independent of the number of physical qubits used for encoding, $N$, while the denominator correctly reflects the cost of the logical operation set. This clearly demonstrates how RCC, by its very definition, automatically strips away the ‘free’ encoding structure and only prices the non-trivial logical information. In Appendix C, we provide more detailed implementation guidance by treating this example alongside another core physical scenario (fixed particle number sector) as canonical case studies.It is noteworthy that, from this global perspective, all pure states ($S(\rho)=0$) within the subspace $H_{R}$ share the same maximal leading term $\log(d_{R})/\log(\Gamma_{R})$. This degeneracy is a direct consequence of the information-theoretic nature of RCC. We will revisit this profound feature in Sec. 6.3, where we demonstrate how this degeneracy is lifted by introducing a physically motivated “observational window,” thereby distinguishing the rich internal structures of different pure states.
2.4 A Note on Units: The Structon
@@
To facilitate the understanding of the physical role of the “information bandwidth” $\log\Gamma_{R}$ in Eq. (2.3) and to provide a unified language for comparing complexity across hardware platforms, this paper introduces an auxiliary unit of measurement specific to RCC. The form of Eq. (2.3), $C_{R}(\rho)=D(\rho||\sigma_{R})/\log\Gamma_{R}$, essentially defines a complexity benchmarked against the alphabet size $\Gamma_{R}$ of a single atomic instruction. To make this physical meaning explicit and to simplify reporting, we define the structon, denoted as $\mathrm{st}_{R}$, as the information-theoretic value corresponding to
@@
$C_{R}(\rho)\ [\mathrm{st}_{R}]=\frac{D_{\text{bits}}(\rho||\sigma_{R})}{\log_{2}\Gamma_{R}}=\frac{D_{\text{nats}}(\rho||\sigma_{R})}{\ln\Gamma_{R}}$ (2.5)
The physical interpretation of this unit is: against the background of a given “structured vacuum” $\sigma_{R}$, 1 $\mathrm{st}_{R}$ represents the complexity required to generate a non-trivial structural unit equivalent to the information of a $\Gamma_{R}$-alphabet”. As an independent and self-consistent notational convention, the introduction of $\mathrm{st}_{R}$ does not alter the form or proof of any theorem in this paper; all correction terms in the main theorem, Theorem 3.1 (spectral skew and logarithmic overhead), when measured in bits or nats, are naturally converted to the $\mathrm{st}_{R}$ unit via Eq. (2.5). To maintain compatibility with traditional information-theoretic expressions, the main text will primarily use the dimensionless form $D/\log\Gamma_{R}$; readers may directly interpret this as a physical quantity in units of $\mathrm{st}_{R}$.A minimal example: Consider a platform with $\Gamma_{R}=12$. If an experimental evaluation yields a relative entropy $D(\rho\|\sigma_{R})=\ln(144)$, the leading term of the complexity lower bound is calculated as $CR(\rho)=D_{\mathrm{nats}}/\ln(\Gamma_{R})=\ln(144)/\ln(12)=2~\mathrm{st}_{R}$. Ignoring logarithmic and spectral corrections, one can report a lower bound on the optimal circuit complexity as $C_{\mathrm{opt}}(\rho)\gtrsim 2~\mathrm{st}_{R}$. A standardized report for the leading term itself would be: $CR(\rho)=2~\mathrm{st}_{R}\;[\Gamma_{R}=12]$. To ensure clarity and reproducibility, we recommend following a minimal set of guidelines when reporting results using $\mathrm{st}_{R}$:
@@
(Note: This section, Sec. 2.4, only provides the reporting and conversion rules for $\mathrm{st}_{R}$; its internal logarithmic base convention does not supersede the conventions in other parts of this paper. The logarithmic base throughout this paper defaults to the established convention of each respective section and appendix (such as Sec. 2.1 and Appendix A); this section has explicitly specified $\log_{2}$ and $\ln$ in Eq. (2.4) and Eq. (2.5) to ensure no ambiguity. As a reporting unit, $\mathrm{st}_{R}$ does not change the form of any theorem or proof.)
3 Main Results: A Structurally Fair Lower Bound for Circuit Complexity

### v3

@@ Abstract
I Introduction
@@
Quantifying the intrinsic complexity of many-body quantum states is a central problem at the interface between computational theory and fundamental physics, with implications for characterizing the resource requirements of quantum computation, the black hole information paradox [Hawking1976, Page1993] and its later developments [AMPS2013], the emergence of spacetime geometry [Maldacena1997, RyuTakayanagi2006], and even the characterization of quantum phases of matter. At the core of answering this question lies the quantification of the minimum computational resources required to prepare a specific quantum state. This perspective is not only key to assessing the potential of quantum algorithms but also provides a fundamental benchmark for measuring the true performance of quantum hardware, with concrete diagnostics developed in recent works [JeffersonMyers2017, Chapman2018]. Although various profound paradigms for complexity measures have been developed in the field—including Nielsen’s complexity based on geometric shortest paths[Nielsen2006], Krylov complexity based on operator algebra growth [parker2019universal], and resource theories closely related to quantum computational advantage (e.g., magic-state and related approaches [Veitch2014, Howard2014], recent links to operator growth [craps2024relation], and general resource-theoretic formalisms [BrandaoGour2015])—a long-standing challenge remains to be systematically addressed: how to conduct a fair complexity accounting for a “structured” quantum system. The so-called “structural fairness problem” originates from the ubiquitous presence of encodings and symmetries in quantum information processing, which can dramatically reshape what should count as a genuine computational cost [Nielsen2006]. Taking a fault-tolerant quantum computer as an example, its information is encoded within a certain “code subspace” of a vast physical Hilbert space; for a basic operation executed on a logical qubit, its true computational cost should depend on the change in the logical information, rather than the apparent consumption of mobilizing many physical qubits to maintain the error-correcting code structure [Gottesman1997, Fowler2012]. In practice, logical operations may be realized via nontrivial transformations supported by the code architecture itself [BravyiKitaev2005, GidneyFowler2019]. If a complexity measure cannot distinguish between these two, it could lead to the misleading conclusion that the larger the physical scale of the system and the more complex the encoding structure, the higher its complexity is assessed to be [BrandaoGour2015]. Similarly, in symmetry-protected topological (SPT) phases or strongly correlated systems, the low-energy physics is confined to specific symmetry sectors; a fair measure ought to treat these structures—endowed “for free” by the Hamiltonian’s symmetries—as an a priori background, and only measure the cost of generating non-trivial quantum correlations on top of this background [ChenGuWen2011, ChenGuLiuWen2013]. This viewpoint is consistent with broader perspectives on emergent phases and constrained dynamics [Senthil2015]. Existing frameworks for measures often find it difficult to systematically isolate these “free structures,” leading to a lack of a unified and fair benchmark when making cross-comparisons of complexity across different physical systems or encoding schemes [Veitch2014, Howard2014]. To directly address this challenge, this paper establishes a theory for quantum complexity lower bounds that is structurally fair, theoretically rigorous, and operationally accessible. Its core is the introduction and systematic construction of the theoretical framework of “Reference-Contingent Complexity” (RCC), and its core contributions can be summarized into the following three points:
@@
The structure of this paper is organized as follows. We first systematically establish the theoretical framework for Reference-Contingent Complexity (RCC) in Sec. II, and then proceed in Sec. III to propose and prove the central theorem of its role as a strict lower bound for circuit complexity. To connect this theory with physical reality, Sec. IV delves into the operational principles of RCC. Building on this foundation, Sec. V presents the axiomatic properties of RCC and its three major application interfaces, while Sec. VI discusses it within a broader theoretical context. Sec. VII provides a summary of the entire paper. To ensure the self-consistency of the manuscript and the completeness of the arguments, all technical derivations are systematically organized in the appendices. Among them, the core theoretical cornerstones—the rigorous proof of the main theorem, the robustness of the framework, and the operational protocols—are elaborated in detail in Appendices A, B, C, and D, respectively. Building upon this, we further extend the boundaries of the theory: Appendices E and F explore the connections of this framework with other theories and its generalization to broader settings, while Appendix G provides constructive evidence for the central conjecture in the main text. Finally, to support the key application interfaces in the main text, Appendix I provides a complete technical supplement on its connection to thermodynamics. A quick reference list of core symbols and terminology can be found in Appendix H.
II Theoretical Framework: Reference-Contingent Complexity (RCC)
@@
II.1 System Setting and Notational Conventions
We consider a finite-dimensional quantum system defined on a Hilbert space $H$ of total dimension $d$. The dynamical evolution of the system is driven by a universal finite gate set $G\subseteq\mathrm{SU}(d)$ with cardinality $g$. Our fundamental objective is to quantify the minimum resource cost required to prepare a target state $\rho$ from a simple initial state (such as a product state) using a quantum circuit composed of this gate set. This minimum resource cost, the optimal circuit complexity $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$, is defined as the length of the shortest sequence of atomic instructions capable of preparing $\rho$ within a precision $\epsilon$ (as measured by the trace distance). To ensure that our complexity benchmark can be compared with concrete physical implementations, we must objectively and unambiguously calibrate the fundamental unit of computational operations. This leads to the core postulate of our framework:
@@ II.1 System Setting and Notational Conventions
The value of $\Gamma_{R}$ is based on a rigorously formalized computational model that combines the gate set $g$ with the hardware’s addressability $|S_{R}|$. A complete exposition of this model, including the precise definition of $\Gamma_{R}=g\cdot|S_{R}|$ and the handling of parallel circuits and continuous-parameter gates, is axiomatically defined in detail in Appendix A. This postulate directly links abstract computational steps with the physical capabilities of the hardware (gate set size and addressing range), enabling our theory to capture the essential differences in computational power among different physical platforms. To ensure the self-consistency of the theory, we list here the key information-theoretic quantities that will be used in the subsequent derivations. The von Neumann entropy of a quantum state $\rho$ is defined as $S(\rho)=-\mathrm{Tr}(\rho\cdot\log(\rho))$[HiaiPetz1991]. The quantum relative entropy between two quantum states $\rho$ and $\sigma$ is $D(\rho||\sigma)=\mathrm{Tr}[\rho\cdot(\log(\rho)-\log(\sigma))]$;[NielsenChuang2010] on the other hand, throughout this paper, to maintain consistency with information-theoretic conventions, all logarithms are assumed to be base-2 (units in bits) by default. In Sec. V.4 and Appendix I, where we interface with the framework of thermodynamics, we will primarily use the natural logarithm (units in nats) for convenience of calculation, and will make explicit conversions when necessary.
@@
II.2 Reference Set and Structured Priors
@@
II.3 Definition of Reference-Contingent Complexity
@@
$C_{R}(\rho)=\frac{1}{\log\Gamma_{R}}D(\rho||\sigma_{R})$ (2)
@@
Eq. (3) reveals the physical meaning of RCC. Its numerator term, $(\log d_{R}-S(\rho))$, represents the “information deficit” or “degree of order” of the quantum state $\rho$ within its respective structured subspace $H_{R}$. $\log d_{R}$ is the maximum Shannon information capacity of this subspace (i.e., $S(\sigma_{R})$), while $S(\rho)$ is the von Neumann entropy actually contained in the state $\rho$. The difference between the two is precisely the degree to which the state deviates from “complete disorder,” and it is also precisely the amount of non-trivial information that must be “injected” by any quantum circuit that prepares the state. At the same time, the denominator term $\log\Gamma_{R}$ provides a precise calibration of the fundamental unit of computational resources. Therefore, the physical picture of RCC is: the structurally fair complexity of a quantum state is equal to the minimum amount of information required to be injected for its preparation, divided by the maximum information bandwidth that a single complete computational operation can provide. To specifically illustrate how RCC achieves structural fairness, we once again consider a stabilizer code that encodes $k$ logical qubits into $N$ physical qubits. The dimension of its code subspace $H_{R}$ is $d_{R}=2^{k}$. For a logical pure state $\rho_{L}$ within the code space, its von Neumann entropy is $S(\rho_{L})=0$. Assuming logical gates can act arbitrarily on the $k$ logical qubits with a universal gate set of cardinality $g_{L}$, and that a single operation has an addressable range covering all $k$ logical qubits, i.e., $|S_{R}|=k$. According to Eq. (3) and the definition $\Gamma_{R}=g_{L}\cdot|S_{R}|$, its RCC is $C_{R}(\rho_{L})=k/\log(g_{L}\cdot k)$. This result aligns with physical intuition: the complexity of this logical pure state is proportional to the number of logical information bits it carries, $k$, and is independent of the number of physical qubits used for encoding, $N$, while the denominator correctly reflects the cost of the logical operation set. This clearly demonstrates how RCC, by its very definition, automatically strips away the ‘free’ encoding structure and only prices the non-trivial logical information. In Appendix C, we provide more detailed implementation guidance by treating this example alongside another core physical scenario (fixed particle number sector) as canonical case studies.It is noteworthy that from this global perspective, all pure states ($S(\rho)=0$) within the subspace $\mathcal{H}_{R}$ share the same maximal leading term $\log(d_{R})/\log(\Gamma_{R})$. This degeneracy is a direct consequence of the information-theoretic nature of RCC. We will revisit this profound feature in Sec. VI.5, where we demonstrate how this degeneracy is lifted by introducing a physically motivated “observational window,” thereby distinguishing the rich internal structures of different pure states.
II.4 A Note on Units: The Structon
@@
To facilitate the understanding of the physical role of the “information bandwidth” $\log\Gamma_{R}$ in Eq. (3) and to provide a unified language for comparing complexity across hardware platforms, this paper introduces an auxiliary unit of measurement specific to RCC. The form of Eq. (3), $C_{R}(\rho)=D(\rho||\sigma_{R})/\log\Gamma_{R}$, essentially defines a complexity benchmarked against the alphabet size $\Gamma_{R}$ of a single atomic instruction. To make this physical meaning explicit and to simplify reporting, we define the structon, denoted as $\mathrm{st}_{R}$, as the information-theoretic value corresponding to
@@
$C_{R}(\rho)\ [\mathrm{st}_{R}]=\frac{D_{\text{bits}}(\rho||\sigma_{R})}{\log_{2}\Gamma_{R}}=\frac{D_{\text{nats}}(\rho||\sigma_{R})}{\ln\Gamma_{R}}$ (5)
The physical interpretation of this unit is: against the background of a given “structured vacuum” $\sigma_{R}$, 1 $\mathrm{st}_{R}$ represents the complexity required to generate a non-trivial structural unit equivalent to the information of a $\Gamma_{R}$-alphabet”. As an independent and self-consistent notational convention, the introduction of $\mathrm{st}_{R}$ does not alter the form or proof of any theorem in this paper; all correction terms in the main theorem, Theorem 3.1 (spectral skew and logarithmic overhead), when measured in bits or nats, are naturally converted to the $\mathrm{st}_{R}$ unit via Eq. (5). To maintain compatibility with traditional information-theoretic expressions, the main text will primarily use the dimensionless form $D/\log\Gamma_{R}$; readers may directly interpret this as a physical quantity in units of $\mathrm{st}_{R}$.A minimal example: Consider a platform with $\Gamma_{R}=12$. If an experimental evaluation yields a relative entropy $D(\rho\|\sigma_{R})=\ln(144)$, the leading term of the complexity lower bound is calculated as $CR(\rho)=D_{\mathrm{nats}}/\ln(\Gamma_{R})=\ln(144)/\ln(12)=2~\mathrm{st}_{R}$. Ignoring logarithmic and spectral corrections, one can report a lower bound on the optimal circuit complexity as $C_{\mathrm{opt}}(\rho)\gtrsim 2~\mathrm{st}_{R}$. A standardized report for the leading term itself would be: $CR(\rho)=2~\mathrm{st}_{R}\;[\Gamma_{R}=12]$. To ensure clarity and reproducibility, we recommend following a minimal set of guidelines when reporting results using $\mathrm{st}_{R}$:
@@
(Note: This section, Sec. II.4, only provides the reporting and conversion rules for $\mathrm{st}_{R}$; its internal logarithmic base convention does not supersede the conventions in other parts of this paper. The logarithmic base throughout this paper defaults to the established convention of each respective section and appendix (such as Sec. II.1 and Appendix A); this section has explicitly specified $\log_{2}$ and $\ln$ in Eq. (4) and Eq. (5) to ensure no ambiguity. As a reporting unit, $\mathrm{st}_{R}$ does not change the form of any theorem or proof.)
III Main Results: A Structurally Fair Lower Bound for Circuit Complexity